OpenStudy (anonymous):

Write the equation of a circle with a center at (-5, -7) and a radius of 3.8

OpenStudy (anonymous):

Equation of circle with center at \((h.k)\) and radius \(r\) is given by: \[(x - h)^2 + (y-k)^2 = r^2\] Here h = -5, k = -7 and r = 3.8 Just plug in the values and find the equation..

OpenStudy (anonymous):

(x--5)^2+(y--7)^2=3.8^2 ?

OpenStudy (anonymous):

@waterineyes ??

OpenStudy (anonymous):

Yes.. Remember: - - = +

OpenStudy (anonymous):

So can you simplify now ??

OpenStudy (anonymous):

Just replace -- by +..

OpenStudy (anonymous):

k...thanks!

OpenStudy (anonymous):

Welcome dear..